Which term of the arithmetic sequence is -18 , given a1=7 and a2=2?
Found 2 solutions by Aldorozos, ikleyn:
Answer by Aldorozos(172)

Read here
http://www.regentsprep.org/regents/math/algtrig/atp2/arithseq.htm
Your first term is 7
The second term is
2=7+(2-1)r
r=-5
-18= 7+ (n-1)(-5)
N=6

Answer by ikleyn(52788)

.
Or read here
    - Arithmetic progressions
    - The proofs of the formulas for arithmetic progressions
    - Problems on arithmetic progressions
    - Word problems on arithmetic progressions
    - Chocolate bars and arithmetic progressions
    - Free fall and arithmetic progressions
    - Uniformly accelerated motions and arithmetic progressions
    - Increments of a quadratic function form an arithmetic progression
    - Mathematical induction and arithmetic progressions
    - One characteristic property of arithmetic progressions
    - Mathematical induction for sequences other than arithmetic or geometric
    - Proving inequalities by the method of Mathematical Induction
    - Solved problems on arithmetic progressions

OVERVIEW of my lessons on arithmetic progressions with short annotations is in the lesson  OVERVIEW of lessons on arithmetic progressions.